\footnotesize
\begin{thebibliography}{99}

\bibitem{MZ}
J.~A.~Minahan and K.~Zarembo,
%``The Bethe-ansatz for N = 4 super Yang-Mills,''
JHEP {\bf 0303} (2003) 013
[arXiv:hep-th/0212208].

\bibitem{gauge}
N.~Beisert, C.~Kristjansen and M.~Staudacher,
%``The dilatation operator of N = 4 super Yang-Mills theory,''
Nucl.\ Phys.\ B {\bf 664} (2003) 131
[arXiv:hep-th/0303060].
$\bullet$
N.~Beisert,
%``The complete one-loop dilatation operator of N = 4 super Yang-Mills
%theory,''
Nucl.\ Phys.\ B {\bf 676} (2004) 3
[arXiv:hep-th/0307015].
$\bullet$
N.~Beisert and M.~Staudacher,
%``The N = 4 SYM integrable super spin chain,''
Nucl.\ Phys.\ B {\bf 670} (2003) 439
[arXiv:hep-th/0307042].
$\bullet$
N.~Beisert,
%``The su(2$|$3) dynamic spin chain,''
Nucl.\ Phys.\ B {\bf 682} (2004) 487
[arXiv:hep-th/0310252].
$\bullet$
N.~Beisert, V.~Dippel and M.~Staudacher,
%``A novel long range spin chain and planar N = 4 super Yang-Mills,''
JHEP {\bf 0407} (2004) 075
[arXiv:hep-th/0405001].
$\bullet$
M.~Staudacher,
%``The factorized S-matrix of CFT/AdS,''
JHEP {\bf 0505} (2005) 054
[arXiv:hep-th/0412188].
$\bullet$
N.~Beisert and M.~Staudacher,
%``Long-range PSU(2,2$|$4) Bethe ansaetze for gauge theory and strings,''
Nucl.\ Phys.\ B {\bf 727} (2005) 1
[arXiv:hep-th/0504190].

\bibitem{string}
I.~Bena, J.~Polchinski and R.~Roiban,
%``Hidden symmetries of the AdS(5) x S**5 superstring,''
Phys. \ Rev. \ D {\bf 69} (2004) 046002
[arXiv:hep-th/0305116].
$\bullet$
G.~Arutyunov, S.~Frolov, J.~Russo and A.~A.~Tseytlin,
%``Spinning strings in AdS(5) x S**5 and integrable systems,''
Nucl.\ Phys.\ B {\bf 671} (2003) 3
[arXiv:hep-th/0307191].
$\bullet$
G.~Arutyunov, J.~Russo and A.~A.~Tseytlin,
%``Spinning strings in AdS(5) x S**5: New integrable system relations,''
Phys.\ Rev.\ D {\bf 69} (2004) 086009
[arXiv:hep-th/0311004].
$\bullet$
V.~A.~Kazakov, A.~Marshakov, J.~A.~Minahan and K.~Zarembo,
%``Classical / quantum integrability in AdS/CFT,''
JHEP {\bf 0405} (2004) 024
[arXiv:hep-th/0402207].
$\bullet$
G.~Arutyunov, S.~Frolov and M.~Staudacher,
%``Bethe ansatz for quantum strings,''
JHEP {\bf 0410} (2004) 016
[arXiv:hep-th/0406256].
$\bullet$
G.~Arutyunov and S.~Frolov,
%``Integrable Hamiltonian for classical strings on AdS(5) x S**5,''
JHEP {\bf 0502} (2005) 059
[arXiv:hep-th/0411089].
$\bullet$
L.~F.~Alday, G.~Arutyunov and A.~A.~Tseytlin,
%``On integrability of classical superstrings in AdS(5) x S**5,''
JHEP {\bf 0507} (2005) 002
[arXiv:hep-th/0502240].
$\bullet$
%\cite{Minahan:2005jq}
  J.~A.~Minahan,
  %``The SU(2) sector in AdS/CFT,''
  Fortsch.\ Phys.\  {\bf 53}, 828 (2005)
  [arXiv:hep-th/0503143].
  %%CITATION = FPYKA,53,828;%%
$\bullet$
S.~Frolov, J.~Plefka and M.~Zamaklar,
%``The AdS(5) x S**5 superstring in light-cone gauge and its Bethe
%equations,''
J.\ Phys.\ A  {\bf 39} (2006) 13037
[arXiv:hep-th/0603008].
%%CITATION = JPAGB,A39,13037;%%
$\bullet$
G.~Arutyunov and S.~Frolov,
%``On AdS(5) x S**5 string S-matrix,''
Phys.\ Lett.\ B {\bf 639} (2006) 378
[arXiv:hep-th/0604043].
$\bullet$
D.~M.~Hofman and J.~M.~Maldacena,
%``Giant magnons,''
J.\ Phys.\ A  {\bf 39} (2006) 13095
[arXiv:hep-th/0604135].
$\bullet$
%\cite{Dorey:2006dq}
  N.~Dorey,
  %``Magnon bound states and the AdS/CFT correspondence,''
  J.\ Phys.\ A  {\bf 39}, 13119 (2006)
  [arXiv:hep-th/0604175].
  %%CITATION = JPAGB,A39,13119;%%
$\bullet$
%\cite{Maldacena:2006rv}
  J.~M.~Maldacena and I.~Swanson,
  %``Connecting giant magnons to the pp-wave: An interpolating limit of   $AdS_5
  %\times S^5$,''
  Phys.\ Rev.\  D {\bf 76}, 026002 (2007)
  [arXiv:hep-th/0612079].
  %%CITATION = PHRVA,D76,026002;%%
$\bullet$
%\cite{Klose:2007rz}
  T.~Klose, T.~McLoughlin, J.~A.~Minahan and K.~Zarembo,
  %``World-sheet scattering in AdS(5) x S**5 at two loops,''
  JHEP {\bf 0708}, 051 (2007)
  [arXiv:0704.3891 [hep-th]].
  %%CITATION = JHEPA,0708,051;%%
$\bullet$
%\cite{Puletti:2007hq}
  V.~Giangreco Marotta Puletti, T.~Klose and O.~Ohlsson Sax,
  %``Factorized world-sheet scattering in near-flat AdS_5 x S^5,''
  Nucl.\ Phys.\  B {\bf 792}, 228 (2008)
  [arXiv:0707.2082 [hep-th]].
  %%CITATION = NUPHA,B792,228;%%
$\bullet$
%\cite{Zarembo:2008hb}
  K.~Zarembo,
  %``Quantum Giant Magnons,''
  JHEP {\bf 0805}, 047 (2008)
  [arXiv:0802.3681 [hep-th]].
  %%CITATION = JHEPA,0805,047;%%
$\bullet$
%\cite{Dorey:2008zy}
  N.~Dorey,
  ``A Spin Chain from String Theory,''
  arXiv:0805.4387 [hep-th].
  %%CITATION = ARXIV:0805.4387;%%

\bibitem{bl}
S.~Frolov, J.~Plefka and M.~Zamaklar,
%``The AdS(5) x S**5 superstring in light-cone gauge and its Bethe
%equations,''
J.\ Phys.\ A  {\bf 39} (2006) 13037
[arXiv:hep-th/0603008].
%%CITATION = JPAGB,A39,13037;%%
$\bullet$
T.~Klose, T.~McLoughlin, R.~Roiban and K.~Zarembo,
%``Worldsheet scattering in AdS(5) x S**5,''
JHEP {\bf 0703} (2007) 094
[arXiv:hep-th/0611169].

\bibitem{rev}
%\cite{Tseytlin:2003ii}
  A.~A.~Tseytlin,
  ``Spinning strings and AdS/CFT duality,''
  arXiv:hep-th/0311139.
  %%CITATION = HEP-TH/0311139;%%
$\bullet$
N.~Beisert,
%``The dilatation operator of N = 4 super Yang-Mills theory and
%integrability,''
Phys.\ Rept.\  {\bf 405} (2005) 1
[arXiv:hep-th/0407277].
$\bullet$
  J.~Plefka,
  %``Spinning strings and integrable spin chains in the AdS/CFT
  %correspondence,''
  Living Rev.\ Rel.\  {\bf 8}, 9 (2005)
  [arXiv:hep-th/0507136].
  %%CITATION = 00222,8,9;%%
$\bullet$
J.~A.~Minahan,
%``A Brief Introduction To The Bethe Ansatz In N=4 Super-Yang-Mills,''
J.\ Phys.\ A  {\bf 39}, 12657 (2006).
%%CITATION = JPAGB,A39,12657;%%

\bibitem{B}
N.~Beisert,
``The su(2$|$2) dynamic S-matrix,''
arXiv:hep-th/0511082.

\bibitem{nlin}
 N.~Beisert,
%``The Analytic Bethe Ansatz for a Chain with Centrally Extended
%su(2$|$2) Symmetry,''
J.\ Stat.\ Mech.\  {\bf 0701}, P017 (2007)
[arXiv:nlin.si/0610017].
%%CITATION = JSTAT,0701,P017;%%

\bibitem{Gleb}
G.~Arutyunov, S.~Frolov and M.~Zamaklar,
%``The Zamolodchikov-Faddeev algebra for AdS(5) x S**5 superstring,''
JHEP {\bf 0704}, 002 (2007)
[arXiv:hep-th/0612229].
%%CITATION = JHEPA,0704,002;%%

\bibitem{J}
R.~A.~Janik,
%``The AdS(5) x S**5 superstring worldsheet S-matrix and crossing
%symmetry,''
Phys.\ Rev.\ D {\bf 73} (2006) 086006
[arXiv:hep-th/0603038].

\bibitem{GHPST}
C.~Gomez and R.~Hernandez,
%``The magnon kinematics of the AdS/CFT correspondence,''
JHEP {\bf 0611} (2006) 021
[arXiv:hep-th/0608029].
$\bullet$
J.~Plefka, F.~Spill and A.~Torrielli,
%``On the Hopf algebra structure of the AdS/CFT S-matrix,''
Phys.\ Rev.\ D {\bf 74} (2006) 066008
[arXiv:hep-th/0608038].

\bibitem{BYang}
N.~Beisert,
%``The S-Matrix of AdS/CFT and Yangian Symmetry,''
PoS {\bf SOLVAY}, 002 (2006)
[arXiv:0704.0400 [nlin.SI]].
%%CITATION = POSCI,SOLVAY,002;%%

\bibitem{MMT}
%\cite{Matsumoto:2007rh}
  T.~Matsumoto, S.~Moriyama and A.~Torrielli,
  %``A Secret Symmetry of the AdS/CFT S-matrix,''
  JHEP {\bf 0709}, 099 (2007)
  [arXiv:0708.1285 [hep-th]].
  %%CITATION = JHEPA,0709,099;%%  

\bibitem{BS}
  N.~Beisert and F.~Spill,
  ``The Classical r-matrix of AdS/CFT and its Lie Bialgebra Structure,''
  arXiv:0708.1762 [hep-th].
  %%CITATION = ARXIV:0708.1762;%%

\bibitem{finite1}
  G.~Arutyunov and S.~Frolov,
  %``On String S-matrix, Bound States and TBA,''
  JHEP {\bf 0712}, 024 (2007)
  [arXiv:0710.1568 [hep-th]].
  %%CITATION = JHEPA,0712,024;%%

\bibitem{finite2}
G.~Arutyunov, S.~Frolov and M.~Zamaklar,
  %``Finite-size effects from giant magnons,''
  Nucl.\ Phys.\  B {\bf 778}, 1 (2007)
  [arXiv:hep-th/0606126].
  %%CITATION = NUPHA,B778,1;%%
$\bullet$
D.~Astolfi, V.~Forini, G.~Grignani and G.~W.~Semenoff,
  %``Gauge invariant finite size spectrum of the giant magnon,''
  Phys.\ Lett.\  B {\bf 651}, 329 (2007)
  [arXiv:hep-th/0702043].
  %%CITATION = PHLTA,B651,329;%%
$\bullet$
%\cite{Kotikov:2007cy}
  A.~V.~Kotikov, L.~N.~Lipatov, A.~Rej, M.~Staudacher and V.~N.~Velizhanin,
  %``Dressing and Wrapping,''
  J.\ Stat.\ Mech.\  {\bf 0710}, P10003 (2007)
  [arXiv:0704.3586 [hep-th]].
  %%CITATION = JSTAT,0710,P10003;%%
$\bullet$
  R.~A.~Janik and T.~Lukowski,
  %``Wrapping interactions at strong coupling -- the giant magnon,''
  Phys.\ Rev.\  D {\bf 76}, 126008 (2007)
  [arXiv:0708.2208 [hep-th]].
  %%CITATION = PHRVA,D76,126008;%%
$\bullet$
J.~A.~Minahan and O.~O.~Sax,
  ``Finite size effects for giant magnons on physical strings,''
  arXiv:0801.2064 [hep-th].
  %%CITATION = ARXIV:0801.2064;%%
$\bullet$
  N.~Gromov, S.~Schafer-Nameki and P.~Vieira,
  ``Quantum Wrapped Giant Magnon,''
  arXiv:0801.3671 [hep-th].
  %%CITATION = ARXIV:0801.3671;%%
$\bullet$
%\cite{Heller:2008at}
  M.~P.~Heller, R.~A.~Janik and T.~Lukowski,
  ``A new derivation of Luscher F-term and fluctuations around the giant
  magnon,''
  arXiv:0801.4463 [hep-th].
  %%CITATION = ARXIV:0801.4463;%%
$\bullet$
  Y.~Hatsuda and R.~Suzuki,
  ``Finite-Size Effects for Dyonic Giant Magnons,''
  arXiv:0801.0747 [hep-th].
  %%CITATION = ARXIV:0801.0747;%%
$\bullet$
%\cite{Klose:2008rx}
  T.~Klose and T.~McLoughlin,
  ``Interacting finite-size magnons,''
  arXiv:0803.2324 [hep-th].
  %%CITATION = ARXIV:0803.2324;%%
$\bullet$
%\cite{Ramadanovic:2008qd}
  B.~Ramadanovic and G.~W.~Semenoff,
  ``Finite Size Giant Magnon,''
  arXiv:0803.4028 [hep-th].
  %%CITATION = ARXIV:0803.4028;%%
$\bullet$
%\cite{Fioravanti:2008rv}
  D.~Fioravanti, P.~Grinza and M.~Rossi,
  ``Strong coupling for planar ${\cal N}=4$ SYM theory: an all-order result,''
  arXiv:0804.2893 [hep-th].
  %%CITATION = ARXIV:0804.2893;%%
$\bullet$
%\cite{Janik:2008hs}
  R.~A.~Janik and T.~Lukowski,
  ``From nesting to dressing,''
  arXiv:0804.4295 [hep-th].
  %%CITATION = ARXIV:0804.4295;%%

\bibitem{finite3}
%\cite{Beisert:2005mq}
  N.~Beisert, A.~A.~Tseytlin and K.~Zarembo,
  %``Matching quantum strings to quantum spins: One-loop vs. finite-size
  %corrections,''
  Nucl.\ Phys.\  B {\bf 715}, 190 (2005)
  [arXiv:hep-th/0502173].
  %%CITATION = NUPHA,B715,190;%%
$\bullet$
%\cite{Sieg:2005kd}
  C.~Sieg and A.~Torrielli,
  %``Wrapping interactions and the genus expansion of the 2-point function  of
  %composite operators,''
  Nucl.\ Phys.\  B {\bf 723}, 3 (2005)
  [arXiv:hep-th/0505071].
  %%CITATION = NUPHA,B723,3;%%
$\bullet$
%\cite{SchaferNameki:2005tn}
  S.~Schafer-Nameki, M.~Zamaklar and K.~Zarembo,
  %``Quantum corrections to spinning strings in AdS(5) x S**5 and Bethe  ansatz:
  %A comparative study,''
  JHEP {\bf 0509}, 051 (2005)
  [arXiv:hep-th/0507189].
  %%CITATION = JHEPA,0509,051;%%
$\bullet$
%\cite{Beisert:2005cw}
  N.~Beisert and A.~A.~Tseytlin,
  %``On quantum corrections to spinning strings and Bethe equations,''
  Phys.\ Lett.\  B {\bf 629}, 102 (2005)
  [arXiv:hep-th/0509084].
  %%CITATION = PHLTA,B629,102;%%
$\bullet$
%\cite{Ambjorn:2005wa}
  J.~Ambjorn, R.~A.~Janik and C.~Kristjansen,
  %``Wrapping interactions and a new source of corrections to the spin-chain  /
  %string duality,''
  Nucl.\ Phys.\  B {\bf 736}, 288 (2006)
  [arXiv:hep-th/0510171].
  %%CITATION = NUPHA,B736,288;%%
$\bullet$
%\cite{SchaferNameki:2006ey}
  S.~Schafer-Nameki, M.~Zamaklar and K.~Zarembo,
  %``How accurate is the quantum string Bethe ansatz?,''
  JHEP {\bf 0612}, 020 (2006)
  [arXiv:hep-th/0610250].
  %%CITATION = JHEPA,0612,020;%%
$\bullet$
  B.~Eden,
  ``Boxing with Konishi,''
  arXiv:0712.3513 [hep-th].
  %%CITATION = ARXIV:0712.3513;%%
$\bullet$
  F.~Fiamberti, A.~Santambrogio, C.~Sieg and D.~Zanon,
  ``Wrapping at four loops in N=4 SYM,''
  arXiv:0712.3522 [hep-th].
  %%CITATION = ARXIV:0712.3522;%%
$\bullet$
  C.~A.~Keeler and N.~Mann,
  ``Wrapping Interactions and the Konishi Operator,''
  arXiv:0801.1661 [hep-th].
  %%CITATION = ARXIV:0801.1661;%%

\bibitem{Etingof}
P.~Etingof and O.~Schiffman,
``Lectures on Quantum Groups,''
Lectures in Mathematical Physics, International Press, Boston, 1998.
$\bullet$
V.~Chari and A.~Pressley,
``A Guide To Quantum Groups,''
Cambridge, UK: Univ. Press (1994).
$\bullet$
V.~G.~Drinfeld,
%``Quantum groups,''
Proc. of the International Congress of Mathematicians, Berkeley,
1986, American Mathematical Society, 1987, 798.
$\bullet$
P.~Etingof, I.~Frenkel and A.~Kirillov,
``Lectures on representation theory and Knizhnik-Zamolodchikov equations,''
Math. Surveys and Monographs, vol. 58, Amer. Math. Soc., Providence, RI, 1998.

\bibitem{KT}
V.~G.~Drinfeld,
Soviet Math. Dokl. {\bf 36} (1988) 212.
$\bullet$
S.~M.~Khoroshkin and V.~N.~Tolstoy,
%``Universal R-matrix for quantized (super)algebras,''
Commun. Math. Phys. {\bf 141} (1991) 599.
$\bullet$
S.~M.~Khoroshkin and V.~N.~Tolstoy,
``Yangian Double And Rational R Matrix,''
arXiv:hep-th/9406194.
%%CITATION = HEP-TH/9406194;%%
$\bullet$
V.~Stukopin,
``Quantum Double of Yangian of Lie Superalgebra A(m,n) and computation
of Universal R-matrix,''
arXiv:math/0504302.
$\bullet$
V.~Stukopin,
%''Yangians of Lie superalgebras of type A(m, n),''
Funct. Anal. Appl. {\bf 28} (1994) 217.
$\bullet$
J.~ Cai, S.~ Wang, K.~Wu and C.~ Xiong,
%''Universal R-matrix Of The Super Yangian Double DY(gl(1|1)),''
Comm. Theor. Phys. {\bf 29} (1998) 173 [arXiv:q-alg/9709038].
$\bullet$
V.~Stukopin,
%''Yangians of classical lie superalgebras: Basic constructions, quantum double
%and universal R-matrix,'' 
Proceedings of the Institute of Mathematics of NAS of Ukraine
{\bf 50} (2004) 1195.
$\bullet$
J.~Brundan and A.~Kleshchev, 
%''Parabolic presentations of the Yangian Y(gl_n),''
Commun. Math. Phys. {\bf 254} (2005) 191.
$\bullet$
A.~Molev,
``Yangians and Classical Lie Algebras,''
Mathematical Surveys and Monographs {\bf 143},
American Mathematical Society, Providence, RI, 2007.
$\bullet$
L.~Gow,
%''Gauss decomposition of the Yangian Y(gl_{m|n}),''
Commun. Math. Phys. {\bf 276} (2007) 799.
$\bullet$
H.~Yamane, 
%``Universal R-matrices for quantum groups
%associated to simple Lie superalgebras,''
Proc. Japan Acad. Ser. A {\bf 67} (1991) 108.
$\bullet$
H.~Yamane,
%``Quantized enveloping algebras associated with simple Lie superalgebras
%and their universal R-matrices,''
Publ. Res. Math. Inst. Sci. {\bf 30} (1994) 15.
$\bullet$
H.~Yamane,
%``On defining relations of affine Lie superalgebras and affine quantized
%universal enveloping superalgebras,''
Publ. Res. Math. Inst. Sci. {\bf 35} (1999) 321,
{\bf 37} (2001) 615 (Errata).
[arXiv:q-alg/9603015].
$\bullet$
I.~Heckenberger, F.~Spill, A.~Torrielli and H.~Yamane,
``Drinfeld second realization of the quantum affine superalgebras of 
$D^{(1)}(2,1;x)$ via the Weyl groupoid,'' 
arXiv:0705.1071 [math.QA]

\bibitem{Yangian}
G.~Mandal, N.~V.~Suryanarayana and S.~R.~Wadia,
%``Aspects of semiclassical strings in AdS(5),''
Phys.\ Lett.\  B {\bf 543}, 81 (2002)
[arXiv:hep-th/0206103].
%%CITATION = PHLTA,B543,81;%%
$\bullet$
L.~Dolan, C.~R.~Nappi and E.~Witten,
%``A relation between approaches to integrability in superconformal
%Yang-Mills theory,''
JHEP {\bf 0310}, 017 (2003)
[arXiv:hep-th/0308089].
$\bullet$
L.~Dolan, C.~R.~Nappi and E.~Witten,
``Yangian symmetry in D = 4 superconformal Yang-Mills theory,''
arXiv:hep-th/0401243.
$\bullet$
M.~Hatsuda and K.~Yoshida,
%``Classical integrability and super Yangian of superstring on AdS(5) x
%S**5,''
Adv.\ Theor.\ Math.\ Phys.\  {\bf 9}, 703 (2005)
[arXiv:hep-th/0407044].
%%CITATION = 00203,9,703;%%
$\bullet$
A.~Das, J.~Maharana, A.~Melikyan and M.~Sato,
%``The algebra of transition matrices for the AdS(5) x S**5
%superstring,''
JHEP {\bf 0412}, 055 (2004)
[arXiv:hep-th/0411200].
%%CITATION = JHEPA,0412,055;%%
$\bullet$
A.~Das, A.~Melikyan and M.~Sato,
%``The algebra of flat currents for the string on AdS(5) x S**5 in the
%light-cone gauge,''
JHEP {\bf 0511}, 015 (2005)
[arXiv:hep-th/0508183].
%%CITATION = JHEPA,0511,015;%%
$\bullet$
A.~Agarwal,
``Comments on higher loop integrability in the su(1$|$1) sector of N = 4
SYM: Lessons from the su(2) sector,''
arXiv:hep-th/0506095.
%%CITATION = HEP-TH/0506095;%%
$\bullet$
B.~I.~Zwiebel,
%``Yangian symmetry at two-loops for the su(2$|$1) sector of N = 4 SYM,''
J.\ Phys.\ A  {\bf 40}, 1141 (2007)
[arXiv:hep-th/0610283].
$\bullet$
  N.~Beisert and B.~I.~Zwiebel,
  %``On Symmetry Enhancement in the psu(1,1|2) Sector of N=4 SYM,''
  JHEP {\bf 0710}, 031 (2007)
  [arXiv:0707.1031 [hep-th]].
  %%CITATION = JHEPA,0710,031;%%
$\bullet$
  S.~Aoyama,
  ``Classical Exchange Algebra of the Superstring on S5 with the AdS-time,''
  arXiv:0709.3911 [hep-th].
  %%CITATION = ARXIV:0709.3911;%% 
$\bullet$
N.~Beisert and D.~Erkal,
  %``Yangian Symmetry of Long-Range gl(N) Integrable Spin Chains,''
  J.\ Stat.\ Mech.\  {\bf 0803}, P03001 (2008)
  [arXiv:0711.4813 [hep-th]].
  %%CITATION = JSTAT,0803,P03001;%%
$\bullet$
  A.~Mikhailov and S.~Schaefer-Nameki,
  ``Algebra of transfer-matrices and Yang-Baxter equations on the string
  worldsheet in AdS(5) x S(5),''
  arXiv:0712.4278 [hep-th].
  %%CITATION = ARXIV:0712.4278;%% 
$\bullet$
%\cite{Ihry:2008gm}
  J.~N.~Ihry,
  %``Yangians in Deformed Super Yang-Mills Theories,''
  JHEP {\bf 0804}, 051 (2008)
  [arXiv:0802.3644 [hep-th]].
  %%CITATION = JHEPA,0804,051;%%

%\cite{Kazakov:2007fy}
\bibitem{Kazakov}
  V.~Kazakov, A.~Sorin and A.~Zabrodin,
  %``Supersymmetric Bethe ansatz and Baxter equations from discrete Hirota
  %dynamics,''
  Nucl.\ Phys.\  B {\bf 790}, 345 (2008)
  [arXiv:hep-th/0703147].
  %%CITATION = NUPHA,B790,345;%%
$\bullet$
%\cite{Kazakov:2007na}
  V.~Kazakov and P.~Vieira,
  ``From Characters to Quantum (Super)Spin Chains via Fusion,''
  arXiv:0711.2470 [hep-th].
  %%CITATION = ARXIV:0711.2470;%%

\bibitem{Hubbard}
  A.~Rej, D.~Serban and M.~Staudacher,
  %``Planar N = 4 gauge theory and the Hubbard model,''
  JHEP {\bf 0603}, 018 (2006)
  [arXiv:hep-th/0512077].
  %%CITATION = JHEPA,0603,018;%%
$\bullet$
G.~Feverati, D.~Fioravanti, P.~Grinza and M.~Rossi,
  %``On the finite size corrections of anti-ferromagnetic anomalous  dimensions
  %in N = 4 SYM,''
  JHEP {\bf 0605}, 068 (2006)
  [arXiv:hep-th/0602189].
  %%CITATION = JHEPA,0605,068;%%
$\bullet$
  B.~Eden and M.~Staudacher,
  %``Integrability and transcendentality,''
  J.\ Stat.\ Mech.\  {\bf 0611}, P014 (2006)
  [arXiv:hep-th/0603157].
  %%CITATION = JSTAT,0611,P014;%%
$\bullet$
G.~Feverati, D.~Fioravanti, P.~Grinza and M.~Rossi,
  %``Hubbard's adventures in N = 4 SYM-land? Some non-perturbative
  %considerations on finite length operators,''
  J.\ Stat.\ Mech.\  {\bf 0702}, P001 (2007)
  [arXiv:hep-th/0611186].
  %%CITATION = JSTAT,0702,P001;%%
$\bullet$
C.~Gomez and R.~Hernandez,
%``Quantum deformed magnon kinematics,''
JHEP {\bf 0703}, 108 (2007)
[arXiv:hep-th/0701200].
%%CITATION = JHEPA,0703,108;%%
$\bullet$
%\cite{Martins:2007hb}
  M.~J.~Martins and C.~S.~Melo,
  %``The Bethe ansatz approach for factorizable centrally extended S
  %-matrices,''
  Nucl.\ Phys.\  B {\bf 785}, 246 (2007)
  [arXiv:hep-th/0703086].
  %%CITATION = NUPHA,B785,246;%%
$\bullet$
%\cite{Stefanski:2007dp}
  B.~J.~Stefanski,
  %``Landau-Lifshitz sigma-models, fermions and the AdS/CFT correspondence,''
  JHEP {\bf 0707}, 009 (2007)
  [arXiv:0704.1460 [hep-th]].
  %%CITATION = JHEPA,0707,009;%%
$\bullet$
%\cite{de Leeuw:2007uf}
  M.~de Leeuw,
  %``Coordinate Bethe Ansatz for the String S-Matrix,''
  J.\ Phys.\ A  {\bf 40}, 14413 (2007)
  [arXiv:0705.2369 [hep-th]].
  %%CITATION = JPAGB,A40,14413;%%
$\bullet$
%\cite{Young:2007wd}
  C.~A.~S.~Young,
  %``q-Deformed Supersymmetry and Dynamic Magnon Representations,''
  J.\ Phys.\ A  {\bf 40}, 9165 (2007)
  [arXiv:0704.2069 [hep-th]].
  %%CITATION = JPAGB,A40,9165;%%
$\bullet$
  D.~Fioravanti and M.~Rossi,
  %``On the commuting charges for the highest dimension SU(2) operator in planar
  %${\cal N}=4$ SYM,''
  JHEP {\bf 0708}, 089 (2007)
  [arXiv:0706.3936 [hep-th]].
  %%CITATION = JHEPA,0708,089;%%
$\bullet$
%\cite{Mansson:2007jg}
  T.~Mansson,
  %``Is there a tower of charges to be discovered?,''
  J.\ Phys.\ A  {\bf 41}, 194014 (2008)
  [arXiv:0711.0931 [hep-th]].
  %%CITATION = JPAGB,A41,194014;%%
$\bullet$
%\cite{Gomez:2007st}
  C.~Gomez, J.~Gunnesson and R.~Hernandez,
  ``The Ising model and planar N=4 Yang-Mills,''
  arXiv:0711.3404 [hep-th].
  %%CITATION = ARXIV:0711.3404;%%

%\cite{Beisert:2008tw}
\bibitem{qu}
  N.~Beisert and P.~Koroteev,
  %``Quantum Deformations of the One-Dimensional Hubbard Model,''
  J.\ Phys.\ A  {\bf 41}, 255204 (2008)
  [arXiv:0802.0777 [hep-th]].
  %%CITATION = JPAGB,A41,255204;%%

\bibitem{Moriyama}
  T.~Matsumoto and S.~Moriyama,
  %``An Exceptional Algebraic Origin of the AdS/CFT Yangian Symmetry,''
  JHEP {\bf 0804}, 022 (2008)
  [arXiv:0803.1212 [hep-th]].
  %%CITATION = JHEPA,0804,022;%%

\bibitem{T}
A.~Torrielli,
%``Classical r-matrix of the su(2$|$2) SYM spin-chain,''
Phys. Rev. D {\bf 75} (2007) 105020
[arXiv:hep-th/0701281].

\bibitem{MT}
  S.~Moriyama and A.~Torrielli,
  %``A Yangian Double for the AdS/CFT Classical r-matrix,''
  JHEP {\bf 0706}, 083 (2007)
  [arXiv:0706.0884 [hep-th]].
  %%CITATION = JHEPA,0706,083;%%

%\cite{deLeeuw:2008dp}
\bibitem{Leeuw}
  M.~de Leeuw,
  ``Bound States, Yangian Symmetry and Classical r-matrix for the AdS5 x S5
  Superstring,''
  arXiv:0804.1047 [hep-th].
  %%CITATION = ARXIV:0804.1047;%%

%\cite{Arutyunov:2008zt}
\bibitem{GS}
  G.~Arutyunov and S.~Frolov,
  ``The S-matrix of String Bound States,''
  arXiv:0803.4323 [hep-th].
  %%CITATION = ARXIV:0803.4323;%%

%\cite{Spill:2008tp}
\bibitem{FA}
  F.~Spill and A.~Torrielli,
  ``On Drinfeld's second realization of the AdS/CFT $su(2|2)$ Yangian,''
  arXiv:0803.3194 [hep-th].
  %%CITATION = ARXIV:0803.3194;%%

\bibitem{BHLBES}
N.~Beisert, R.~Hernandez and E.~Lopez,
%``A crossing-symmetric phase for AdS(5) x S**5 strings,''
JHEP {\bf 0611} (2006) 070
[arXiv:hep-th/0609044].
%%CITATION = PHRVA,D75,085010;%%
$\bullet$
N.~Beisert, B.~Eden and M.~Staudacher,
%``Transcendentality and crossing,''
J.\ Stat.\ Mech.\  {\bf 0701}, P021 (2007)
[arXiv:hep-th/0610251].
%%CITATION = JSTAT,0701,P021;%%

\bibitem{BCDKS}
Z.~Bern, M.~Czakon, L.~J.~Dixon, D.~A.~Kosower and V.~A.~Smirnov,
%``The Four-Loop Planar Amplitude and Cusp Anomalous Dimension in
%Maximally Supersymmetric Yang-Mills Theory,''
Phys.\ Rev.\  D {\bf 75}, 085010 (2007)
[arXiv:hep-th/0610248].
%%CITATION = PHRVA,D75,085010;%%
$\bullet$
M.~K.~Benna, S.~Benvenuti, I.~R.~Klebanov and A.~Scardicchio,
%``A test of the AdS/CFT correspondence using high-spin operators,''
Phys.\ Rev.\ Lett.\ {\bf 98}, 131603 (2007)
[arXiv:hep-th/0611135].
%%CITATION = PRLTA,98,131603;%%
$\bullet$
L.~F.~Alday, G.~Arutyunov, M.~K.~Benna, B.~Eden and I.~R.~Klebanov,
%``On the strong coupling scaling dimension of high spin operators,''
JHEP {\bf 0704}, 082 (2007)
[arXiv:hep-th/0702028].
%%CITATION = JHEPA,0704,082;%%
$\bullet$
M.~Beccaria, G.~F.~De Angelis and V.~Forini,
%``The scaling function at strong coupling from the quantum string Bethe 
%equations,''
JHEP {\bf 0704}, 066 (2007)
[arXiv:hep-th/0703131].
%%CITATION = JHEPA,0704,066;%%
$\bullet$
P.~Y.~Casteill and C.~Kristjansen,
%``The Strong Coupling Limit of the Scaling Function from the Quantum   String
%Bethe Ansatz,''
Nucl.\ Phys.\  B {\bf 785}, 1 (2007)
[arXiv:0705.0890 [hep-th]].
%%CITATION = NUPHA,B785,1;%%

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage{amsmath,amssymb,hyperref,epsfig,multicol}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}

\newcommand{\beq}{\begin{equation}}
\newcommand{\eeq}{\end{equation}}
\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\eea}{\end{eqnarray}}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}

\newcommand{\cR}{{\mathcal R}}

\newcommand{\fL}{{\mathfrak L}}
\newcommand{\fR}{{\mathfrak R}}
\newcommand{\fQ}{{\mathfrak Q}}
\newcommand{\fS}{{\mathfrak S}}

\newcommand{\alg}[1]{\mathfrak{#1}}

\usepackage{epsfig,multicol}
\begin{document}

\begin{titlepage}
\begin{flushright} 
MIT-CTP 3953\\
\end{flushright}
\mbox{ }  \hfill 
\vspace{5ex}
\Large
\begin {center}     
{\bf Structure of the string R-matrix}
\end {center}
\large
\vspace{1ex}
\begin{center}
Alessandro Torrielli
\end{center}
\vspace{1ex}
\begin{center}
Center for Theoretical Physics\\
Laboratory for Nuclear Sciences\\
and\\
Department of Physics\\
Massachusetts Institute of Technology\\
Cambridge, Massachusetts 02139, USA\\ 
\vspace{1ex}
\texttt{torriell@mit.edu}

\end{center}
\vspace{4ex}
\rm
\begin{center}
{\bf Abstract}
\end{center} 
\normalsize 
By requiring invariance directly under the Yangian symmetry, we rederive
Beisert's quantum R-matrix, in a form that carries explicit dependence on the 
representation labels, the braiding factors, and the spectral parameters 
$u_i$. In this way, we demonstrate that there exist a rewriting of its entries,
such that
the dependence on the spectral parameters is purely {\it of difference form}.
Namely, the latter enter only in the combination $u_1-u_2$, as indicated by the shift automorphism of the Yangian. When 
recasted in this fashion, the entries exhibit a cleaner structure, which 
allows to 
spot new interesting relations among them. This permits to package them into a 
practical tensorial expression, where the non-diagonal entries are taken care 
by explicit combinations of symmetry algebra generators.

\vfill
\end{titlepage} 

\section{Introduction}
Integrability in AdS/CFT \cite{MZ,gauge,string,bl} (see \cite{rev} for reviews)
is strictly related to the existence of an
R-matrix \cite{B}, 
which is a solution of the quantum Yang-Baxter equation \cite{nlin,Gleb}, and 
satisfies the Hopf-algebraic analog of crossing symmetry \cite{J,GHPST}.  
This R-matrix exhibits a certain infinite-dimensional Yangian-type symmetry 
\cite{BYang},
based on the central extension of the Lie superalgebra $\alg{psu(2|2)}$,
plus an additional Yangian generator of $\alg{u(2|2)}$ signature 
\cite{MMT,BS}. A universal form of the R-matrix would be desirable, 
particularly in view of its potential use in studying finite-size effects 
\cite{finite1,finite2,finite3}.  
Such universal R-matrix is still unknown, mainly due to the 
fact that the dual Coxeter number of the algebra vanishes, and 
therefore traditional mathematical techniques do not 
straightforwardly apply \cite{Etingof,KT}. 
Nevertheless, huge progress has come in several 
aspects \cite{Yangian,Kazakov,Hubbard,qu,Moriyama}, like in the study of 
the correspondent classical r-matrix \cite{T,MT,BS,Moriyama,Leeuw}, 
in deriving higher representations \cite{GS}, and in giving the Yangian
an almost canonical form \cite{FA}.

The way the R-matrix was originally derived \cite{B} was by using the Lie
superalgebra symmetry in the fundamental representation. 
The combination of the peculiar features of the tensor products of two 
fundamentals \cite{nlin}, 
together with the non-trivial braiding of the coproduct 
\cite{GHPST,bl,Gleb}
allowed to fix all entries up to an 
overall scalar factor \cite{BHLBES,BCDKS}. Yangian symmetry was 
explicitly discovered 
only later \cite{BYang}, 
initially hardly visible due to the consistency 
relations that mix together all the 
parameters in the game. In higher
representations the Yangian 
may impose useful conditions \cite{Leeuw}, but the 
Yang-Baxter equation could probably be sufficient to fix the R-matrix 
\cite{GS}.
 
Were it not for the algebraic peculiarity mentioned above, 
then probably, as it happens for more traditional Yangians,
the Lie superalgebra symmetry 
would not be enough, and the invariance under the 
Yangian would determine much of the form of the R-matrix.
Since the Yangian carries the spectral parameter $u$, this normally 
gives the familiar
rational R-matrices, depending only on the difference $u_1 - u_2$ due to a constant
shift automorphism $u \rightarrow u+c$. 
From there, investigation of the universal R-matrix is easier. 

In this paper, we want to exploit this privileged viewpoint that the Yangian
gives on the R-matrix, and imagine we were to fix the latter by purely using 
the (first level) Yangian generators. Our strategy is to try to avoid explicit
use of the constraints, and to carry on the various parameters almost until
the end as if they were unrelated. 
In particular, the relevant Yangian still possesses the shift automorphism \cite{BYang,FA}.
Our calculation produces explicit expressions for the entries, where the dependence on the spectral 
parameters $u_i$ is indeed purely {\it of difference form}, namely $u_1 - u_2$, as expected for the above reasons. Moreover, the other parameters (representation labels, braiding factors) are also recognizable, which helps in finding new relations among
the entries, and tells us much about the combination of generators 
that is going to 
produce them. Upon imposing the constraints, we 
exactly recover the original quantum R-matrix found by Beisert.
Tempted by this new perspective, 
we also provide a tensorial repackaging of the 
entries which further clarifies their hidden structure, and might serve as
a device to investigate the properties of the universal R-matrix.

\section{Structural relations}
In this section, we will follow the approach of rederiving the quantum 
R-matrix, by 
using the 
fundamental representation not of the Lie superalgebra, as it is traditionally 
done, but rather using from the very beginning the first Yangian generators.
The correspondent 
coproducts are given in \cite{BYang,FA}. In doing this, we will
leave explicit the dependence on the representation labels, the braiding 
factors, and the spectral parameters, as if they were not related to each 
other. This will allow us to trace them back 
through the entries of the R-matrix.
Of course, only when they satisfy the familiar 
relations one can find a solution,
therefore it is impossible to really 
terminate the process without imposing those 
relations. Nevertheless, it is possible to proceed further enough to 
reach a satisfactory 
form for all the entries, which we report below. At that point, there is no 
need to continue, and we can just read the entries we obtain. 
As a check, we can then verify that, on the constraints, we recover
Beisert's quantum R-matrix.

Let us first remind the basic definition of the labels of the
fundamental representation: 

\begin{align}
&\fR^a{}_b|\phi^c\rangle=\delta^c_b|\phi^a\rangle - \frac{1}{2} \delta^a_b|\phi^c\rangle~,\quad 
\fL^\alpha{}_\beta|\phi^\gamma\rangle=\delta^\gamma_\beta|\phi^\alpha\rangle - \frac{1}{2} \delta^\alpha_\beta|\phi^\gamma\rangle~,\nonumber\\
&\fQ^\alpha{}_a|\phi^b\rangle= a \, \delta^b_a|\psi^\alpha\rangle~,\quad
\fQ^\alpha{}_a|\psi^\beta\rangle
= b \, \epsilon^{\alpha\beta}\epsilon_{ab}|\phi^b\rangle~,\nonumber\\
&\fS^a{}_\alpha|\phi^b\rangle
= c \, \epsilon^{ab}\epsilon_{\alpha\beta}|\psi^\beta\rangle~,\quad
\fS^a{}_\alpha|\psi^\beta\rangle= d \, \delta^\beta_\alpha|\phi^a\rangle~.
\end{align}
The labels satisfy $ad-bc=1$.
The central (braiding) operator $\alg{U}$ has eigenvalue $U$, and $u$ will be 
the spectral parameter.

Let us write the R-matrix in the familiar $\alg{su(2)}\oplus \alg{su(2)}$ 
invariant fashion: 

\begin{align}
\cR_{12}|\phi^a_1\phi^b_2\rangle
&=R^{12}_{12}|\phi^a_1\phi^b_2\rangle
+R^{12}_{21}|\phi^b_1\phi^a_2\rangle
+R^{12}_{34} \, \epsilon^{ab}\epsilon_{\alpha\beta}
|\psi^\alpha_1\psi^\beta_2\rangle~,\nonumber\\
\cR_{12}|\psi^\alpha_1\psi^\beta_2\rangle
&=R^{34}_{34}|\psi^\alpha_1\psi^\beta_2\rangle
+R^{34}_{43}|\psi^\beta_1\psi^\alpha_2\rangle
+R^{34}_{12} \, \epsilon^{\alpha\beta}\epsilon_{ab}
|\phi^a_1\phi^b_2\rangle~,\nonumber\\
\cR_{12}|\phi^a_1\psi^\beta_2\rangle
&=R^{13}_{13}|\phi^a_1\psi^\beta_2\rangle
+R^{13}_{31}\psi^\beta_1\phi^a_2\rangle~,\nonumber\\
\cR_{12}|\psi^\alpha_1\phi^b_2\rangle
&=R^{31}_{31}|\psi^\alpha_1\phi^b_2\rangle
+R^{31}_{13}|\phi^b_1\psi^\alpha_2\rangle~.
\end{align}
Choosing the overall normalization to be such that $R^{33}_{33} = R^{34}_{34} + R^{34}_{43}  \, = \, U_2/U_1$,  
the expressions for the entries which we obtain by imposing invariance under the 
Yangian are given 
by the following
formulas:

\begin{align}
\label{a}
&R^{31}_{31} = U_2 \, \frac{u_1 - u_2}{u_1 - u_2 - a_1 d_1 + 
a_1 b_1 (c_2/a_2) \, U_2^2}~,\nonumber\\
&R^{31}_{13} = R^{31}_{31} \, \frac{1}{U_2} \, \frac{(a_2 d_1 - 
b_1 c_2 \, U_2^2)}{u_1 - u_2}~,\nonumber\\
&R^{13}_{13} = \frac{U_2^2}{U_1} \, \frac{u_1 - u_2}{u_1 - u_2 + b_2 
c_2 - a_2 b_2 (d_1/b_1) \, U_1^2}~,\nonumber\\
&R^{13}_{31} = R^{13}_{13} \, \frac{1}{U_2} \, \frac{(- b_2 c_1 + 
a_1 d_2 \, U_2^2)}{u_1 - u_2}~,\nonumber\\
&R^{12}_{34}= R^{12}_{12} \, \frac{1}{U_2} \, \frac{(- c_1 a_2 + 
a_1 c_2 \, U_2^2)}{u_1 - u_2 - 1}~,\nonumber\\ 
&R^{12}_{21} = R^{12}_{12} \, \frac{1}{u_1 - u_2} \, \bigg( 1 + \frac{1}{U_2^2} \, \frac{(c_1 a_2 - 
a_1 c_2 \, U_2^2)(d_1 b_2 - 
b_1 d_2 \, U_2^2)}{u_1 - u_2 - 1}\bigg)~,\nonumber\\
&R^{34}_{12}= R^{34}_{34} \, \frac{1}{U_2} \, \frac{(d_1 b_2 - 
b_1 d_2 \, U_2^2)}{u_1 - u_2 + 1}~,\nonumber\\
&R^{34}_{43} = R^{34}_{34} \, \frac{1}{u_1 - u_2} \, \bigg(- 1 + \frac{1}{U_2^2}\, \frac{(c_1 a_2 - 
a_1 c_2 \, U_2^2)(d_1 b_2 - 
b_1 d_2 \, U_2^2)}{u_1 - u_2 + 1}\bigg)~, 
\end{align}

\begin{align}
\label{b}
&R^{11}_{11}= R^{12}_{12} +R^{12}_{21} = \frac{U_2^2}{U_1^2} \, \frac{u_1 - u_2 - b_1 c_1 + 
a_1 b_1 (d_2/b_2) \, U_2^2}{u_1 - u_2 + b_2 c_2 - 
a_2 b_2 (d_1/b_1) \, U_1^2}~,\\ 
&\nonumber
\end{align}

As one can immediately see, the dependence of the entries on the
spectral parameters $u_i$ is purely {\it of difference form}, as indicated by 
the fact that the Yangian is invariant under an automorphism that shifts the
spectral parameter by a constant. 

One can check that these expressions indeed reproduce the entries of Beisert's R-matrix \cite{B}. It is sufficient to recall the parametrization 
\begin{align}
\label{c}
a=\sqrt{g}\gamma~,\quad
b=\sqrt{g}\frac{\alpha}{\gamma}\biggl(1-\frac{x^+}{x^-}\biggr)~,\quad
c=\sqrt{g}\frac{i\gamma}{\alpha x^+}~,\quad
d=\sqrt{g}\frac{x^+}{i\gamma}\biggl(1-\frac{x^-}{x^+}\biggr)~,
\end{align}
together with $U=\sqrt{x^+/x^-}$ and the relation
$x^++1/x^+-x^--1/x^-=i/g$, and that $u = (ig/2)\, (x^+ + x^-)(1 + 1/x^+ x^-)$.
Then, we compare with the familiar formulas

\begin{align}
\cR_{12}|\phi^a_1\phi^b_2\rangle
&=\frac{1}{2}(A_{12}-B_{12})|\phi^a_1\phi^b_2\rangle
+\frac{1}{2}(A_{12}+B_{12})|\phi^b_1\phi^a_2\rangle
+\frac{1}{2}C_{12}\epsilon^{ab}\epsilon_{\alpha\beta}
|\psi^\alpha_1\psi^\beta_2\rangle~,\nonumber\\
\cR_{12}|\psi^\alpha_1\psi^\beta_2\rangle
&=-\frac{1}{2}(D_{12}-E_{12})|\psi^\alpha_1\psi^\beta_2\rangle
-\frac{1}{2}(D_{12}+E_{12})|\psi^\beta_1\psi^\alpha_2\rangle
-\frac{1}{2}F_{12}\epsilon^{\alpha\beta}\epsilon_{ab}
|\phi^a_1\phi^b_2\rangle~,\nonumber\\
\cR_{12}|\phi^a_1\psi^\beta_2\rangle
&=G_{12}|\phi^a_1\psi^\beta_2\rangle
+H_{12}|\psi^\beta_1\phi^a_2\rangle~,\nonumber\\
\cR_{12}|\psi^\alpha_1\phi^b_2\rangle
&=L_{12}|\psi^\alpha_1\phi^b_2\rangle
+K_{12}|\phi^b_1\psi^\alpha_2\rangle~,
\end{align}
with the functions $A_{12},B_{12},\ldots$ given by

\begin{align}
\label{d}
&A_{12}=\frac{x_2^+-x_1^-}{x_2^--x_1^+}~,\quad
B_{12}=\frac{x_2^+-x_1^-}{x_2^--x_1^+}
\biggl(1-2\frac{1-1/x_1^+x_2^-}{1-1/x_1^+x_2^+}
\frac{x_2^--x_1^-}{x_2^+-x_1^-}\biggr)~,\nonumber\\
&\qquad C_{12}=\frac{2\gamma_1\gamma_2U_2}{\alpha x_1^+x_2^+}
\frac{1}{1-1/x_1^+x_2^+}
\frac{x_2^--x_1^-}{x_2^--x_1^+}~,\nonumber\\
&D_{12}=-\frac{U_2}{U_1}~,\quad
E_{12}=-\frac{U_2}{U_1}
\biggl(1-2\frac{1-1/x_1^-x_2^+}{1-1/x_1^-x_2^-}
\frac{x_2^+-x_1^+}{x_2^--x_1^+}\biggr)~,\nonumber\\
&\qquad F_{12}=-\frac{2\alpha(x_1^+-x_1^-)(x_2^+-x_2^-)}
{\gamma_1\gamma_2U_1x_1^-x_2^-}
\frac{1}{1-1/x_1^-x_2^-}
\frac{x_2^+-x_1^+}{x_2^--x_1^+}~,\nonumber\\
&G_{12}=\frac{1}{U_1}\frac{x_2^+-x_1^+}{x_2^--x_1^+}~,\quad
H_{12}=\frac{\gamma_1U_2}{\gamma_2U_1}
\frac{x_2^+-x_2^-}{x_2^--x_1^+}~,\nonumber\\
&L_{12}=U_2\frac{x_2^--x_1^-}{x_2^--x_1^+}~,\quad
K_{12}=\frac{\gamma_2}{\gamma_1}
\frac{x_1^+-x_1^-}{x_2^--x_1^+}~.
\end{align}
By making use of (\ref{a}), (\ref{b}), (\ref{c}), (\ref{d}), it is possible to verify that the two 
R-matrices indeed exactly match.

As an interesting remark, we notice that in formulas (\ref{a}) and (\ref{b})
it is possible to {\it formally} switch off the braiding ($U_i\to 1$), 
and also
send everywhere  
$b_i$, $c_i$ to zero\footnote{Undetermined expressions like $b_1/b_2$ are 
sent to $1$.}, and $a_i$, $d_i$ to $1$. 
This should correspond to scatter two representations
of $\alg{gl(2|2)}$. Indeed, if one does that, one finds out that the R-matrix 
becomes formally equal to

\begin{equation}
\cR_{12}^{lim} = \frac{u_1 - u_2}{u_1 - u_2 - 1} \, \bigg( 1\otimes 1 
+ \frac{1}{u_1 - u_2} \sum_{i,j=1}^4 \, (-)^j \, E_{ij} \otimes E_{ji}\bigg)~,
\end{equation}
where $E_{ij}$ are the unit matrices with all zeroes but $1$ in position 
$(i,j)$, and bosonic and fermionic indices are altogether 
numbered from $1$ to $4$. The combination 
$\sum_{i,j=1}^4 \, (-)^j \, E_{ij} \otimes E_{ji}$ is the quadratic 
Casimir of $\alg{gl(2|2)}\otimes \alg{gl(2|2)}$, which already emerged from one-loop gauge theory \cite{gauge,finite1},
and from the classical
analysis of \cite{T}. The R-matrix which we obtain by this formal procedure is 
recognized this time as the 
quantum (Yang-type) R-matrix of $\alg{gl(2|2)}$ in the fundamental 
representation, and it is well known to solve the 
Yang-Baxter equation. We have thus find a consistent practical way of tuning 
on and off 
the central extensions
(which are proportional to $b$ and $c$ respectively) 
in the formula for the R-matrix.

\section{Tensorial repackaging}
We would like to exploit here the rewriting achieved in the previous section,
in order to express the whole R-matrix in a more compact tensorial form.
This will be far from enough to be able to
provide the universal R-matrix, but it will be
a useful exercise which may teach us where some of the terms are likely to
come from.
We (re)write here below 
some of the important identities that one obtains from the 
expressions given in the previous section:

\begin{align}
&R^{31}_{13} = R^{31}_{31} \, \frac{1}{U_2} \, \frac{(a_2 d_1 - 
b_1 c_2 \, U_2^2)}{u_1 - u_2}~,\nonumber\\
&R^{13}_{31} = R^{13}_{13} \, \frac{1}{U_2} \, \frac{(- b_2 c_1 + 
a_1 d_2 \, U_2^2)}{u_1 - u_2}~,\nonumber\\
&R^{12}_{34}= R^{12}_{12} \, \frac{1}{U_2} \, \frac{(- c_1 a_2 + 
a_1 c_2 \, U_2^2)}{u_1 - u_2} \, \frac{u_1 - u_2}{u_1 - u_2 - 1}~,\nonumber\\ 
&R^{34}_{12}= R^{34}_{34} \, \frac{1}{U_2} \, \frac{(d_1 b_2 - 
b_1 d_2 \, U_2^2)}{u_1 - u_2} \, \frac{u_1 - u_2}{u_1 - u_2 + 1}~,\nonumber\\
&R^{12}_{21} = R^{12}_{12} \, \frac{1}{U_2^2} \, \frac{(c_1 a_2 - 
a_1 c_2 \, U_2^2)}{u_1 - u_2} \, \frac{(d_1 b_2 - 
b_1 d_2 \, U_2^2)}{u_1 - u_2} \, \frac{u_1 - u_2}{u_1 - u_2 -1} + R^{12}_{12} \, \frac{1}{u_1 - u_2}~,\nonumber\\
&R^{34}_{43} = R^{34}_{34} \, \frac{1}{U_2^2} \, \frac{(c_1 a_2 - 
a_1 c_2 \, U_2^2)}{u_1 - u_2} \, \frac{(d_1 b_2 - 
b_1 d_2 \, U_2^2)}{u_1 - u_2} \, \frac{u_1 - u_2}{u_1 - u_2 +1} - R^{34}_{34} \, \frac{1}{u_1 - u_2}~.
\end{align}
This seems to suggest a not too unfamiliar 
matrix pattern. After trying different 
possibilities, one can see that a relatively simple 
tensorial repackaging of the R-matrix 
that is able to produce these relations can be found:

\begin{align}
&\cR_{12}= \cR^F_1 \, \cR^F_2 \, \cR^H_{12} + \cR^B_1 \, \cR^H_1 + \cR^B_2 \, \cR^H_2 ~,    
\end{align}
where 
\begin{align}
&\cR^F_1 = 1 \otimes 1 + \frac{1}{u_1 - u_2} \, (\fQ^1{}_1 \otimes \fS^1{}_1 \alg{U} + 
\fQ^2{}_2 \otimes \fS^2{}_2 \alg{U} - \fS^1{}_1 \otimes \fQ^1{}_1 \alg{U}^{-1} - 
\fS^2{}_2 \otimes \fQ^2{}_2 \alg{U}^{-1})~,\nonumber\\
&\cR^F_2 = 1 \otimes 1 + \frac{1}{u_1 - u_2} \, (\fQ^1{}_2 \otimes \fS^2{}_1 \alg{U} + 
\fQ^2{}_1 \otimes \fS^1{}_2 \alg{U} - \fS^1{}_2 \otimes \fQ^2{}_1 \alg{U}^{-1} - 
\fS^2{}_1 \otimes \fQ^1{}_2 \alg{U}^{-1})~,\nonumber\\
&\nonumber\\
&\cR^B_1 = \frac{\Pi_B \otimes \Pi_B}{1 - u_1 + u_2} + \frac{1}{u_1 - u_2} \, (\fR^1{}_2 \otimes \fR^2{}_1 + 
\fR^2{}_1 \otimes \fR^1{}_2)~, \quad\quad \quad \cR^H_1 = R^{12}_{12} \, \, 1\otimes 1~,\nonumber\\
&\cR^B_2 = \frac{\Pi_F \otimes \Pi_F}{1 + u_1 - u_2} - \frac{1}{u_1 - u_2} \, (\fL^1{}_2 \otimes \fL^2{}_1 + 
\fL^2{}_1 \otimes \fL^1{}_2)~, \quad\quad \quad \, \, \, \, \cR^H_2 = R^{34}_{34} \, \, 1\otimes 1~,\nonumber\\
&\nonumber\\
&\cR^H_{12} = R^{13}_{13} \, \, \Pi_B \otimes \Pi_F \, + R^{31}_{31} \, \, \Pi_F \otimes \Pi_B \, +  \cR^H_B~,\\ \nonumber
\end{align}
$\Pi_B$ and $\Pi_F$ being the projectors onto the 
bosonic and fermionic subspaces respectively, $\Pi_B = diag\{1,1,0,0\}$ and 
$\Pi_F = diag\{0,0,1,1\}$. 
The only non-zero entries of $\cR^H_B$ are
\begin{align}
&\cR^H_B |\phi^a_1 \phi^a_2 \rangle
=\bigg( R^{11}_{11} + \frac{R^{12}_{12}}{u_1 - u_2 - 1} \bigg)\, |\phi^a_1 \phi^a_2 \rangle~,\quad \, \, \, \, 
\cR^H_B \, \epsilon_{a b} |\phi^a_1 \phi^b_2 \rangle= R^{12}_{12} \, \frac{u_1 - u_2}{u_1 - u_2 - 1} \, \epsilon_{a b} |\phi^a_1 \phi^b_2 \rangle~,\nonumber\\
&R^H_B |\psi^\alpha_1 \psi^\alpha_2 \rangle
=\bigg( R^{33}_{33} - \frac{R^{34}_{34}}{u_1 - u_2 + 1} \bigg)\, |\psi^\alpha_1 \psi^\alpha_2 \rangle~,\quad
\cR^H_B \, \epsilon_{\alpha \beta} |\psi^\alpha_1 \psi^\beta_2 \rangle= R^{34}_{34} \, \frac{u_1 - u_2}{u_1 - u_2 + 1} \,\epsilon_{\alpha \beta } |\psi^\alpha_1 \psi^\beta_2 \rangle~.
\end{align}

This has to be seen as a practical tool to organize the entries, useful for a 
subsequent attempt to find the universal R-matrix. The Cartan part is 
traditionally the most complicated to reproduce, as shown by 
formulas (\ref{a}), (\ref{b}).
But also the root part seems to be a little different from standard cases, 
even when taking into account that in the fundamental representation the roots 
are 
nilpotent, and therefore what appears to be 
linear could well be the first term of a series expansion. 
Nevertheless, in the classical limit,
the formula seems to reproduce the non-diagonal part of the 
classical r-matrix. 
In any case, no conclusions are to be drawn at the moment on the 
persistence of 
this kind of pattern at the universal level, and the 
device is purely practical.

\section{Conclusions}
In this paper we have shown that there exist a way of rewriting Beisert's 
quantum 
R-matrix, such that the dependence on the spectral parameters is purely of 
difference form. This was hidden before by the complicated relations 
connecting the spectral parameter with the representation labels and the braiding factors, and can be
achieved by using the Yangian symmetry. In particular, this dependence
is expected from the shift 
automorphism that the Yangian possesses. The remaining dependence on the 
two representations, which is responsible for the ultimately observed 
non-difference pattern, is encoded in a precise sequence of
representation labels and braiding 
factors. This, in turn, is to be expected, if the R-matrix has to come from some 
universal combination of symmetry generators. We provide some hints at this 
structure, suggested by novel relations among the entries, which in this 
new rewriting are easier to spot.

When comparing with the classical r-matrix analysis of \cite{BS}, one notices
that a suitable readjusting of the classical
variables can be used to abandon a pure $u_1 - u_2$ dependence on the classical
spectral parameters, in order to reach 
a convenient normalization for the extra generator to insert in the classical
Yangian.
The quantum
version of this generator \cite{MMT,BS}
has not yet been canonically embedded in the quantum 
Yangian,
and it is hard to judge on its role at the moment. In other words, it is 
possible that one may 
have to eventually introduced a few uncoupled $u_i$'s in order to 
get a truly universal expression. What we have shown here is that, 
nevertheless, {\it there exist} 
at least one rewriting where the spectral parameters come in 
differences, all the rest being taken care of by recognizable algebraic 
quantities.

\section{Acknowledgments}
The author would like to thank Pavel Etingof and Victor Kac 
for discussion.
Many of the calculations have been performed with \texttt{Mathematica}.
This work is supported in part by funds provided by the U.S. 
Department of Energy (D.O.E.) under cooperative research agreement
DE-FG02-05ER41360. The author thanks Istituto Nazionale di Fisica Nucleare
(I.N.F.N.) for supporting him through a ``Bruno Rossi'' postdoctoral 
fellowship.

\end{document}